For the finite ring R2 = {00000000, 00000001, ..., 11111111} find: 11010111⊖10101010 = N2
Answer #1
  • The finite Ring has elements that are finite in number
  • in other words, the elements are not infinite
  • they are finite and countable with in a set limit
  • The Ring can accommodate binary operations + and . those satisfy the distributive laws
  • For example, for all p,q, and r that belongs to Real set R,
  • we can say that (p+q).r = (p.r + q.r)
  • and also
  • p.(q+r) = p.q + p.r
  • and also
  • q.(p+r) = q.p + q.r
  • which is the same as
  • (p+r).q = p.q + r.q
  • The finite field can make a legitimate example of a finite ring
  • Given a collection of i items, beginning with i = zero, we can make a collection of rings with these i items as follows:
  • a,a,b,b,k,b,d,b,zz,k,d,b,v,b,d,d,cia,b,v,b,v,d,d,b,…..
  • The standard application of finite rings is in terms of the enumerations
  • for each item i that belongs to the ring R, there will be an integer m greater than unity in such a way that i ^ m = i – in that case the ring R satisfies the commutative property
  • as long as there are prime numbers p , and positive integers m, there will be finite fields with a total element count of p^n elements
  • Tow finite elements will satisfy the isomorphic property if their order are the same
  • Hence,
  • 11010111   q 10101010 = N2